OpenStudy (anonymous):

What function best represents a sine function with an amplitude of 4, a period of pi over 6, and a midline at y = −2?

OpenStudy (anonymous):

f(x) = −2 sin(x − pi over 6) + 4 f(x) = 4 sin 12x − 2 f(x) = 4 sin(x − pi over 6) − 2 f(x) = −2 sin 12x + 4

OpenStudy (crashonce):

the amplitude is the coefficient of x

OpenStudy (crashonce):

so 2nd or thid but im not sure sorry

OpenStudy (anonymous):

Thanks, I'm going w/ the third one since it says " a period of pi over 6" in the problem